OpenStudy (anonymous):

5x + 23 = − 2x − 12, which order of operations completely solves for x

OpenStudy (anonymous):

5x + 23 - ( -2x - 12) = 0 ( because they are equal ) 5x + 23 + 2x + 12 = 0 7x + 35 = 0 x = -b/a -35/7 = - 5 5(-5) + 23 = -2(-5) - 12 -25 + 23 = 10 - 12 - 2 = -2